### v1

@@ 6.3 Synergies between intensity and polarization surveys
If we undertake deep integrations that approach the classical intensity confusion limit in the PHI1 band, the PHI2 and PPI band will be limited by confusion in intensity. However, depending on the exact sensitivity ratio between bands, PPI bands may still not be affected by confusion in polarization. In addition, it will also be possible to deblend the PHI2 band using a prior-based extraction algorithm (Donnellan et al., 2024). All the bands will thus be used efficiently in such a strategy, and we will fully exploit the high PPI sensitivity through the polarization. The risk of attempting a first high-z deep polarization survey will also be mitigated, since we will get extremely deep intensity data at shorter wavelength at the same time. PRIMAger is thus a very promising instrument able to open two new windows of survey parameter space with a single deep field observation.
@@ 7 Conclusion
For the projected baseline sensitivities of the PRIMAger wide and deep surveys, the classical confusion limit curve crosses the sensitivity limit at approximately 60 m and 45 m, respectively. Taking advantage of the available instrument sensitivity at longer wavelengths requires using deblending methods. The PRIMAger hyperspectral architecture, which produces finely sampled SEDs, is particularly good at enabling these methods by providing priors for sources detected at shorter, unconfused wavelength. A companion paper (Donnellan et al., 2024), analyzes the performance of a particular deblending approach (XID+, Hurley et al., 2017), showing that its application will recover fluxes out to m and beyond for astrophysical SEDs. Moreover, we showed that in polarization the confusion limit is more than two orders of magnitude lower than in intensity. Surveys will thus be limited by the instrument sensitivity, except at m in the deep field.
@@
We computed the probability to detect a galaxy above the classical confusion limit as a function of its position in the SFR-z plane (Fig. 7 and 10). In intensity, galaxies at the knee of the infrared luminosity function (L) will be above the classical confusion limit in at least one band up to , while massive (10 M) main-sequence galaxies can be recovered up to . In polarization, PRIMager opens up a brand new parameter space by enabling studies for L and massive main-sequence galaxies which are brighter than the classical confusion limit up to and , respectively. We can also observe a tail of extreme objects up to .
We estimate the effect of the background of polarized emission due to high- galaxies on measurements of polarized emission of extended low- foreground sources (Table 3). The 1- noise on polarized surface brightness can vary from 1.2 to 3.4 kJy/sr depending on the band and how bright the foreground is. This noise is correlated between bands, which has an impact on the uncertainties of foreground color measurements. We estimated that a minimum foreground polarized surface brightness of 11 kJy/sr at 96 is necessary to obtain a 10 % precision on the foreground color.
Finally, assuming the PRIMAger baseline sensitivity, we expect several thousands of detections of the integrated dust polarization of high-z galaxies (up to z2.5) in both the deep and wide fields (Fig. 13 and Table 4). Considering that polarization has been reported on only one lensed high-z galaxy so far (Geach et al., 2023), PRIMAger has the potential to revolutionize this field of study by producing large statistical samples.
@@ Acknowledgements.
This research made use of Photutils, an Astropy package fordetection and photometry of astronomical sources (Bradley et al., 2023). James Donnellan was supported by the Science and Technology Facilities Council, grant number ST/W006839/1, through the DISCnet Centre for Doctoral Training. Seb Oliver acknowledges support from the UK Space Agency through ST/Y006038/1. Guilaine Lagache This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 788212). The contributions of the authors using the Contributor Roles Taxonomy (CRediT) were as follows: Matthieu Bethermin: Conceptualization, Methodology, Software, Validation, Investigation, Writing - original Draft; Alberto Bolatto, Laure Ciesla, Jason Glenn, Brandon Hensley, Guilaine Lagache, Enrique Lopez-Rodriguez, Seb Oliver, and Alexandra Pope: Conceptualization, Writing - Review & Editing; Denis Burgarella and James Donnellan: Writing - Review & Editing; Charles Bradford, François Boulanger, and Marc Sauvage: Conceptualization.References
@@ Appendix B Impact of the instrument pixels
In our main analysis, we used a beam FWHM driven only the optics. However, the current design of PRIMAger has pixel sizes that undersample the PSF (particularly at short wavelengths). Pixel sizes range from to for the PHI pixels in the wavelength ranges 23 to 45 m (PHI1) and 45 to 82 m (PHI2). We repeated our analysis using a beam that is broaden by the response of the pixel. To compute this broaden beam, we convolve the diffraction-limited response of the system with the response of the pixel, modeled as a top-hat function with the diameter corresponding to the wavelength range. This estimate is conservative since the effects of pixel size can be mitigated in part by the use of sub-pixel dithers and drizzle-type map making algorithms. These calculations are thus an upper limit on the possible impact. Moreover, the instrument pixel size will likely be further optimized to reduce undersampling effects.

### v3

@@ 6.3 Synergies between intensity and polarization surveys
If we undertake deep integrations that approach the classical intensity confusion limit in the PHI1 band, the PHI2 and PPI band will be limited by confusion in intensity. However, depending on the exact sensitivity ratio between bands, PPI bands may still not be affected by confusion in polarization. In addition, it will also be possible to deblend the PHI2 band using a prior-based extraction algorithm (Donnellan et al., 2024). All the bands will thus be used efficiently in such a strategy, and we will fully exploit the high PPI sensitivity through the polarization. The risk of attempting a first high-z deep polarization survey will also be mitigated, since we will get extremely deep intensity data at shorter wavelength at the same time. PRIMAger is thus a very promising instrument, which is able to open two new windows of survey parameter space with a single deep field observation.
@@ 7 Conclusion
For the conservative estimated sensitivities of the PRIMAger wide and deep surveys, the classical confusion limit curve crosses the sensitivity limit at approximately 60 m and 45 m, respectively. Taking advantage of the available instrument sensitivity at longer wavelengths requires using deblending methods. The PRIMAger hyperspectral architecture, which produces finely sampled SEDs, is particularly good at enabling these methods by providing priors for sources detected at shorter, unconfused wavelength. A companion paper (Donnellan et al., 2024), analyzes the performance of a particular deblending approach (XID+, Hurley et al., 2017), showing that its application will recover fluxes out to m and beyond for astrophysical SEDs. Moreover, we showed that in polarization the confusion limit is more than two orders of magnitude lower than in intensity. Surveys will thus be limited by the instrument sensitivity, except at m in the deep field.
@@
We computed the probability to detect a galaxy above the classical confusion limit as a function of its position in the SFR-z plane (Fig. 7 and 10). In intensity, galaxies at the knee of the infrared luminosity function (L⋆) will be above the classical confusion limit in at least one band up to , while massive (1011 M⊙) main-sequence galaxies can be recovered up to . In polarization, PRIMager opens up a brand new parameter space by enabling studies for L⋆ and massive main-sequence galaxies which are brighter than the classical confusion limit up to and , respectively. We can also observe a tail of extreme objects up to .
We estimate the effect of the background of polarized emission due to high- galaxies on measurements of polarized emission of extended low- foreground sources (Table 3). The 1- noise on polarized surface brightness can vary from 1.2 to 3.4 kJy/sr depending on the band and how bright the foreground is. This noise is correlated between bands, which has an impact on the uncertainties of foreground color measurements. We estimated that a minimum foreground polarized surface brightness of 11 kJy/sr at 96 m is necessary to obtain a 10 % precision on the foreground color.
Finally, assuming the PRIMAger conservative estimated sensitivity, we expect several thousands of detections of the integrated dust polarization of high-z galaxies (up to z2.5) in both the deep and wide fields (Fig. 13 and Table 4). Considering that polarization has been reported on only two lensed high-z galaxy so far (Geach et al., 2023; Chen et al., 2024), PRIMAger has the potential to revolutionize this field of study by producing large statistical samples.
Detecting sources in polarization requires a 100 times better sensitivity than necessary to detect the same sources in intensity at the same wavelength. Opening this new window on high-z dust polarization has thus obviously a cost in observing time. However, it can be offset by the detector intrinsic sensitivity. Indeed, for a deep PRIMAger survey (1 deg2 in 1 500 h) and our conservative estimated sensitivities, both the short-wavelength data in intensity (PHI1) and the polarization maps at longer wavelengths (PPI) will be close to the confusion limit of the instrument. There is thus a strong synergy between pushing the instrument to this limit in intensity at short wavelengths, and performing deep polarization surveys at long wavelengths. At this depth, the long-wavelength data in intensity will be fully dominated by confusion rather than by the extremely low instrumental noise, but analysis methods using short wavelength priors are expected to allow us to deblend sources efficiently using short-wavelength data (Donnellan et al., 2024). A single survey will thus obtain well-sampled mid- to far-infrared SEDs in intensity, opening new opportunities for star formation and AGN science (Bisigello et al., 2024), together with long-wavelength far-infrared data in polarization, providing new constraints on galactic magnetic fields.
@@ Acknowledgements.
This research made use of Photutils, an Astropy package fordetection and photometry of astronomical sources (Bradley et al., 2023). James Donnellan was supported by the Science and Technology Facilities Council, grant number ST/W006839/1, through the DISCnet Centre for Doctoral Training. Seb Oliver acknowledges support from the UK Space Agency through ST/Y006038/1. Guilaine Lagache This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 788212). Portions of this work were carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The contributions of the authors using the Contributor Roles Taxonomy (CRediT) were as follows: Matthieu Bethermin: Conceptualization, Methodology, Software, Validation, Investigation, Writing - original Draft; Alberto Bolatto, Laure Ciesla, Jason Glenn, Brandon Hensley, Guilaine Lagache, Enrique Lopez-Rodriguez, Seb Oliver, and Alexandra Pope: Conceptualization, Writing - Review & Editing; Denis Burgarella and James Donnellan: Writing - Review & Editing; Charles Bradford, François Boulanger, and Marc Sauvage: Conceptualization.References
- André et al. (2019) André, P., Hughes, A., Guillet, V., et al. 2019, PASA, 36, e029
@@
- Chen et al. (2024) Chen, J., Lopez-Rodriguez, E., Ivison, R. J., et al. 2024, arXiv e-prints, arXiv:2407.14596
@@
- Lopez-Rodriguez et al. (2018) Lopez-Rodriguez, E., Alonso-Herrero, A., Diaz-Santos, T., et al. 2018, MNRAS, 478, 2350
@@
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A12
@@ Appendix B Impact of the instrument pixels
In our main analysis, we used a beam FWHM driven only the optics. However, the current design of PRIMAger has pixel sizes that undersample the PSF (particularly at short wavelengths). Pixel sizes range from to for the PHI pixels in the wavelength ranges 23 to 45 m (PHI1) and 45 to 82 m (PHI2). We repeated our analysis using a beam that is broadened by the response of the pixel. To compute this broaden beam, we convolve the diffraction-limited response of the system with the response of the pixel, modeled as a top-hat function with the diameter corresponding to the wavelength range. This estimate is conservative since the effects of pixel size can be mitigated in part by the use of sub-pixel dithers and drizzle-type map making algorithms. These calculations are thus an upper limit on the possible impact. Moreover, the instrument pixel size will likely be further optimized to reduce undersampling effects.
@@
Appendix C Impact of the mean polarization fraction on the expected number of detections
The confusion noise scales linearly with the mean polarization fraction . Since the polarized flux density is also proportional to , the expected number of detections in polarization surveys affected only by confusion is independent of . The situation is more complex in presence of instrumental noise. When decreases, the polarized flux density of the sources and the polarized confusion noise will both decrease, but the instrumental noise will remain constant. In the regime where the instrumental noise is dominant, the number of detections will thus decrease. In this appendix, we discuss a low (=0.7 %) and a high (=1.3 %) mean polarization fraction corresponding to the 3 bounds of the probability density distribution of determined in Appendix 15.
In Table 5 and 6, we present the number of polarized detections expected for a low and high mean polarization fraction, respectively. They can can be compared with the results from our standard simulation (see Table 4). The number of detections remains consistent within a factor of 2 with the standard simulation, except for the wide field in the PPI3 and PPI4 bands assuming payload required sensitivity for which the number of detection is very low in the low case. This uncertainty by a factor of 2 is much smaller than the variations by more than an order of magnitude found for our two different sensitivity scenarios.
Finally, in Fig. 17 and 18, we present the redshift and SFR distributions for our various polarization and sensitivity scenarios. The impact of is mild. As shown in Table 5 and 6, the mean redshift and SFR varies by less than a factor of 2 (after excluding the cases with 10 objects which are not statistically significant). For the conservative estimated sensitivity with larger numbers of detections, the variations are lower than 20 %.
| Band | Plim | Ndet | Mean | Mean SFR |
| Jy | M⊙/yr | |||
| Deep field (1500h, 1deg2) | ||||
| with payload required sensitivity | ||||
| PPI1 | 552 | 46 | 0.22 | 63 |
| PPI2 | 758 | 34 | 0.23 | 75 |
| PPI3 | 1039 | 16 | 0.19 | 72 |
| PPI4 | 1421 | 2 | 0.07 | 11 |
| Wide field (1500h, 10deg2) | ||||
| with payload required sensitivity | ||||
| PPI1 | 1718 | 55 | 0.09 | 18 |
| PPI2 | 2353 | 30 | 0.07 | 11 |
| PPI3 | 3213 | 5 | 0.03 | 2 |
| PPI4 | 4390 | 0 | – | – |
| Deep field (1500h, 1deg2) | ||||
| with conservative estimated sensitivity | ||||
| PPI1 | 61 | 1669 | 0.54 | 66 |
| PPI2 | 105 | 1344 | 0.58 | 84 |
| PPI3 | 178 | 768 | 0.66 | 127 |
| PPI4 | 238 | 352 | 0.86 | 225 |
| Wide field (1500h, 10deg2) | ||||
| with conservative estimated sensitivity | ||||
| PPI1 | 182 | 3145 | 0.40 | 85 |
| PPI2 | 289 | 2345 | 0.43 | 108 |
| PPI3 | 439 | 1155 | 0.47 | 150 |
| PPI4 | 545 | 400 | 0.53 | 212 |
| Band | Plim | Ndet | Mean | Mean SFR |
| Jy | M⊙/yr | |||
| Deep field (1500h, 1deg2) | ||||
| with payload required sensitivity | ||||
| PPI1 | 553 | 136 | 0.32 | 78 |
| PPI2 | 763 | 119 | 0.35 | 96 |
| PPI3 | 1054 | 61 | 0.36 | 128 |
| PPI4 | 1446 | 12 | 0.34 | 177 |
| Wide field (1500h, 10deg2) | ||||
| with payload required sensitivity | ||||
| PPI1 | 1718 | 165 | 0.12 | 20 |
| PPI2 | 2354 | 125 | 0.11 | 17 |
| PPI3 | 3218 | 40 | 0.06 | 9 |
| PPI4 | 4398 | 5 | 0.03 | 2 |
| Deep field (1500h, 1deg2) | ||||
| with conservative estimated sensitivity | ||||
| PPI1 | 69 | 3288 | 0.61 | 57 |
| PPI2 | 135 | 2366 | 0.64 | 75 |
| PPI3 | 255 | 1242 | 0.70 | 113 |
| PPI4 | 362 | 553 | 0.87 | 194 |
| Wide field (1500h, 10deg2) | ||||
| with conservative estimated sensitivity | ||||
| PPI1 | 185 | 8195 | 0.47 | 73 |
| PPI2 | 301 | 6565 | 0.51 | 94 |
| PPI3 | 476 | 3660 | 0.58 | 139 |
| PPI4 | 609 | 1575 | 0.75 | 239 |